OpenStudy (anonymous):

rewite the standard form x-2y=3 into slope-intercept form

OpenStudy (anonymous):

Hi, I'd be happy to help you with this today.

OpenStudy (anonymous):

First, I'd like to know if you know what slope-intercept form is.

OpenStudy (anonymous):

no i really dont

OpenStudy (anonymous):

Slope-intercept form is y = mx + b, where m is the slope of the line and b is the y-intercept.

OpenStudy (anonymous):

What's important to notice is that y is by itself in slope intercept form, and that the x comes first on the right hand side of the =.

OpenStudy (anonymous):

ok thanks :) these teachers dont help me at all

OpenStudy (anonymous):

So, do you have any idea of what the first step would be to get this equation into slope intercept form?

OpenStudy (anonymous):

x-2y=3

OpenStudy (anonymous):

would it be y=x+3

OpenStudy (anonymous):

Wow! You did a lot of work in your head! So, that answer is close to the correct answer but not exactly right.

OpenStudy (anonymous):

Lets retrace your steps together and find the mistake.

OpenStudy (anonymous):

What was your first step?

OpenStudy (anonymous):

well my teacher said to put y=1/2x-3 is that right i dont trust this teacher

OpenStudy (anonymous):

That is also close, but still not quite right.

OpenStudy (anonymous):

The first step in getting the y by itself is to subtract x from both sides.

OpenStudy (anonymous):

x-2y=3 x-2y-x=3-x

OpenStudy (anonymous):

Do you have any questions about this first step?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, any idea of what the next step might be?

OpenStudy (anonymous):

Hint: Simplify the left side of the =.

OpenStudy (anonymous):

no i dont know this at all

OpenStudy (anonymous):

That's okay, I'll help guide you through it.

OpenStudy (anonymous):

On the left side of the = we have an x and a -x. We can combine these two together since they both involve x.

OpenStudy (anonymous):

hey well look this teacher is comming closer to me and wants the answer now can u give it to me please i cant fail

OpenStudy (anonymous):

x - x = 0, so the left side of the equation is -2y.

OpenStudy (anonymous):

Okay, the last step is to divide everything by -2, so -2y = 3 - x becomes y = (1/2)x - 3/2

OpenStudy (anonymous):

Answer: y = (1/2)x - 3/2

OpenStudy (anonymous):

thank u sooo much :)

OpenStudy (anonymous):

You are welcome. Great work and remember to come back to open study and ask more excellent questions!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is there anything else I can help you with today?

OpenStudy (anonymous):

not right now but i will need ur help later on

OpenStudy (anonymous):

Okay, you can become my fan by hovering over my username with your mouse and clicking "Become Fan". Whenever you have a question feel free to ask me or another user for help!

OpenStudy (anonymous):

I have to go now. It was a pleasure working with you today bjones5454. Take care!

OpenStudy (anonymous):

ok :)